OpenStudy (horsegal244):

What is the value of 3^2/3^4? 1/81 1/27 1/9 1/3

OpenStudy (horsegal244):

@seascorpion1

OpenStudy (seascorpion1):

I'm here

OpenStudy (er.mohd.amir):

a^m/a^n=a^(m-n)

OpenStudy (horsegal244):

i dont get that...

OpenStudy (seascorpion1):

3^(2-4)

OpenStudy (er.mohd.amir):

if base is same then power are add here 3^2*3^-4

OpenStudy (seascorpion1):

Yes

OpenStudy (horsegal244):

so i multiply 3 and 3 first right?

OpenStudy (horsegal244):

...

OpenStudy (seascorpion1):

3^2=3*3 3^4=3*3*3*3

OpenStudy (er.mohd.amir):

no add powers tat is 2+(-4)=-2

OpenStudy (seascorpion1):

Just treat it as a fraction

OpenStudy (horsegal244):

3^2=3x3=9 3^4=3x3x3x3=81

OpenStudy (seascorpion1):

9/81=1/9 as 9*9=81

OpenStudy (horsegal244):

ok...

OpenStudy (horsegal244):

these are the answers 1/81 1/27 1/9 1/3

OpenStudy (seascorpion1):

the answer (as I said) is 1/9 (so third option)

OpenStudy (horsegal244):

@seascorpion1

OpenStudy (seascorpion1):

You needing any more help?

OpenStudy (horsegal244):

oh lol ya

OpenStudy (seascorpion1):

where?